# GenExam: A Multidisciplinary Text-to-Image Exam

### v1

@@ Abstract
Exams are a fundamental test of expert-level intelligence and require integrated understanding, reasoning, and generation. Existing exam-style benchmarks mainly focus on understanding and reasoning tasks, and current generation benchmarks emphasize the illustration of world knowledge and visual concepts, neglecting the evaluation of rigorous drawing exams. We introduce GenExam, the first benchmark for multidisciplinary text-to-image exams, featuring 1,000 samples across 10 subjects with exam-style prompts organized under a four-level taxonomy. Each problem is equipped with ground-truth images and fine-grained scoring points to enable a precise evaluation of semantic correctness and visual plausibility. Experiments show that even state-of-the-art models such as GPT-Image-1 and Gemini-2.5-Flash-Image achieve less than 15% strict scores, and most models yield almost 0%, suggesting the great challenge of GenExam. By framing image generation as an exam, GenExam offers a rigorous assessment of models’ ability to integrate knowledge, reasoning, and generation, providing insights on the path to general AGI. Our benchmark and evaluation code is released at https://github.com/OpenGVLab/GenExam.
@@ 3.1 Overview
Taxonomy. In addition to the top-level subject, we also provide a detailed four-level taxonomy, such as “Chemistry/Structure of Matter/Molecular Structure/Organic Compounds” in Fig. 2. The taxonomy annotations are built based on ISCED-F UNESCO (2013) fields, with the aim of improving the academic strictness of GenExam for seamless integration with multidisciplinary research. Detailed four-level taxonomy is provided in Appendix D.
@@ 3.2 Data Curation Pipeline
Image Collection. The data curation pipeline is shown in Fig. 4. To construct a general benchmark for multidisciplinary text-to-image exam, we first consider possible subjects and domains of multidisciplinary images and curate a four-level taxonomy assisted by GPT-5 OpenAI (2025a) and manual check. We then use these taxonomy to generate keywords related to specific subject knowledge, and collect web images based on Google search. Images from existing subject-knowledge-related datasets are also collected: MMMU Yue et al. (2024a), ScienceQA Lu et al. (2022), TextbookQA Kembhavi et al. (2017), MAVIS Zhang et al. (2024), ChEBI Hastings et al. (2016), HLE Phan et al. (2025) and R-Bench Guo et al. (2025b). We obtain approximately 40K images from the data sources. We then conduct automatic filtering by utilizing GPT-5 to rate images in terms of text richness, image domain, complexity and subject knowledge density, and set a threshold to remove undesirable images. 6.5K images are left after automatic filtering. Details of automatic filtering and proportion of image sources are in Appendix B.
@@
Annotating. We employ GPT-5 to generate the initial version of prompts and scoring points. Prompts that require subject knowledge and reasoning are generated, similar to graph-drawing questions in exams. Subsequently, scoring points (questions and scores) are generated based on the prompt. We enforce through prompt engineering that the questions cannot be too general (e.g. “Does the image show the correct structure of benzene?”) or too complex (can be split into multiple sub-questions). Scores are based on the difficulty of the questions, where basic questions should have low scores and harder ones have high scores. We provide some few-shot examples to enhance the labeling process. The complete prompts are provided in Appendix F.2.
@@ 4 Experiments
In this section, we evaluate the performance of representative methods on our GenExam. We also validate the alignment between MLLM-as-a-judge and human evaluation. We use gpt-5-2025-08-07 OpenAI (2025a) as the evaluator model and set the reasoning effort as “low”. More experiments are provided in Appendix C, including results categorized by subject knowledge difficulty and level-2 taxonomy, detailed scores of each dimension, and ablations on the evaluator model.
@@ 4.3 Human Alignment
We conduct a human alignment experiment to examine the validity of MLLM-as-a-judge, following the paradigms in  Zhao et al. (2025); Sun et al. (2025). Five experts are asked to score 250 randomly sampled images generated by GPT-Image-1. They first provide an overall rating between 1 and 10, and then follow the automatic evaluation protocol to annotate scoring points, spelling, logical consistency and readability. The average scores among all experts are used to calculate the mean absolute error (MAE) and correlations between MLLM-judge predicted scores and human scores. In Tab. 3(a), the results show a low MAE across the four dimensions, indicating high precision of our evaluation pipeline. We then use the overall rating to calculate correlations (Kendall’s $\tau$, Spearman’s $\rho$ and Pearson’s $r$) between human scores and four automatic metrics: our relaxed score, our semantic correctness, VQA score Lin et al. (2024), and CLIP score Hessel et al. (2021). In Tab. 3(b), our relaxed score achieves high correlations with human judgments with $p<0.05$, slightly better than using solely semantic correctness without visual plausibility, while CLIP score and VQA score fail to capture the correctness of multidisciplinary images. In Appendix Tab. 4, we further examine the selection of MLLMs as the evaluator model.
4.4 Qualitative Analysis
Examples of generated images are shown in Fig. 7. Taking GPT-Image-1 as an example, common error types include lack of knowledge (e.g. keys in the Circle of Semitones), incorrect reasoning (e.g. failing to infer the coordinate of the intersection point), and graph-drawing errors of logical inconsistency (e.g. the point (1,0)(-1,0) appears to be on the y-axis), highlighting the necessity to integrate multimodal understanding, reasoning and generation. Some models show other graph-drawing errors of spelling (e.g. “Transpostion” for Qwen-Image) or readability (e.g. occluded “Music Theory” for HiDream-I1-Full), which is a key capability in multidisciplinary text-to-image exams. These examples align with our quantitative results: while top closed-source models can partially integrate basic knowledge into generation, they still struggle with precise, domain-aware execution; open-source models, on the contrary, often fail at basic aspects like generating visual plausible images. This underscores the need for future models to prioritize multidisciplinary knowledge integration, rigorous reasoning, and fine-grained visual coherence. Such failures also demonstrate the value of GenExam ’s fine-grained scoring criteria, as they capture flaws that broad metrics like photorealism or general aesthetics would miss. More qualitative results are provided in Appendix G.
5 Conclusion
Ethics Statement
All images in our benchmark are constructed from publicly available resources and existing datasets. We will release the benchmark and the subset with clear usage guidelines to prevent misuse for generating misleading content. We have checked for potential biases across all the subjects, aiming for balanced knowledge coverage without favoring specific demographics or disciplines.
Appendix A More Statistics
(a) Biology
(b) Chemistry
(c) Computer Science
(d) Economics
(e) Engineering
(f) Geography
(g) History
(h) Math
(i) Music
(j) Physics
Figure 10: Word clouds of prompts in each subject.
Appendix B Automatic Data Filtering
Appendix C Additional Experiments
C.1 Ablation on Evaluator Models
@@
We further study the selection of evaluator model by examining different MLLMs, including closed-source models (GPT-5, GPT-5-mini, o3, o4-mini, GPT-4.1, GPT-4o OpenAI (2025a; c)) and open-source models (InternVL3.5, Intern-S1, Intern-S1-mini Wang et al. (2025a); Bai et al. (2025a)). On the same manually evaluated samples as in Tab. 3, we run each of the models as evaluator and calculate the accuracy of answering scoring points, MAE of visual plausibility, and three correlation metrics (Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$). As shown in Tab. 4, the latest model GPT-5 with low reasoning effort achieves the best performance with acceptable cost and time, while using medium or high reasoning effort leads to a potential decrease in performance and higher costs. In addition, using ground truth images as reference help to improve the evaluation precision. Early closed-source models have relatively low performance, especially those without reasoning abilities. For open-source models, even the latest models still observe a large gap between the closed-source ones, where the scoring point accuracy is close to random guess (50%) and some correlations are even of no statistical significance, suggesting metrics with these evaluators are not applicable. This is because multidisciplinary images are still difficult for current multimodal understanding models Yue et al. (2024b). Therefore, it is necessary to use advanced MLLMs to evaluate the generated images for higher precision and robustness.
C.2 Detailed Comparison of Existing Methods
In Tab. 5, we show the results on GenExam-Mini. The performance of the models is consistent with the results from the full dataset, which can be considered as a valid subset for efficient and affordable validation.
@@ C.2 Detailed Comparison of Existing Methods
The results on each level-2 taxonomy is provided in Tab. 8. To highlight the difference between models, and since some models can only obtain a 0% strict score, we only report relaxed scores in Tab. 8.
@@
Model Strict Score Relaxed Score

### v2

@@ Abstract
Exams are a fundamental test of expert-level intelligence and require integrated understanding, reasoning, and generation. Existing exam-style benchmarks mainly focus on understanding and reasoning tasks, and current generation benchmarks emphasize the illustration of world knowledge and visual concepts, neglecting the evaluation of rigorous drawing exams. We introduce GenExam, the first benchmark for multidisciplinary text-to-image exams, featuring 1,000 samples across 10 subjects with exam-style prompts organized under a four-level taxonomy. Each problem is equipped with ground-truth images and fine-grained scoring points to enable a precise evaluation of semantic correctness and visual plausibility. Experiments show that even state-of-the-art models such as GPT-Image-1 and Gemini-2.5-Flash-Image achieve less than 15% strict scores, and most models yield almost 0%, suggesting the great challenge of GenExam. By framing image generation as an exam, GenExam offers a rigorous assessment of models’ ability to integrate understanding, reasoning, and generation, providing insights on the path to general AGI. Our benchmark and evaluation code are released at https://github.com/OpenGVLab/GenExam.
@@ 3.1 Overview
Taxonomy. In addition to the top-level subject, we also provide a detailed four-level taxonomy, such as “Mathematics/Plane Geometry/Circles/Tangent” in Fig. 2. The taxonomy annotations are built based on ISCED-F UNESCO (2013) fields, with the aim of improving the academic strictness of GenExam for seamless integration with multidisciplinary research. Detailed four-level taxonomy is provided in Appendix D.
@@ 3.2 Data Curation Pipeline
Data Collection. The data curation pipeline is shown in Fig. 4. To construct a general benchmark for multidisciplinary text-to-image exam, we first consider possible subjects and domains of multidisciplinary images and curate a four-level taxonomy assisted by GPT-5 OpenAI (2025a) and through manual check. We then use these taxonomy to generate keywords related to specific subject knowledge, and collect web images from exams and textbooks based on Google search. Images from existing subject-knowledge-related datasets are also collected: MMMU Yue et al. (2024a), ScienceQA Lu et al. (2022), TextbookQA Kembhavi et al. (2017), MAVIS Zhang et al. (2024), ChEBI Hastings et al. (2016), HLE Phan et al. (2025) and R-Bench Guo et al. (2025b). We obtain approximately 40K images from the data sources with their original text descriptions (e.g. webpage content for web images, lecture notes in TextbookQA, or question and answers in MMMU). We then conduct automatic filtering by utilizing GPT-5 to rate images in terms of text richness, image domain, complexity and subject knowledge density, and set a threshold to remove undesirable images. 6.5K images are left after automatic filtering. Details of automatic filtering and proportion of image sources are in Appendix B.
@@
Annotating. We employ GPT-5 to generate the initial version of prompts and scoring points. Prompts that require subject knowledge and reasoning are generated, similar to graph-drawing questions in exams. The original text description from the data source is used for reference to improve the precision, bu should be treated carefully since they may contain irrelevant information. Subsequently, scoring points (questions and scores) are generated based on the prompt. We enforce through prompt engineering that the questions cannot be too general (e.g. “Does the image show the correct structure of benzene?”) or too complex (can be split into multiple sub-questions). Scores are based on the difficulty of the questions, where basic questions should have low scores and harder ones have high scores. We provide some few-shot examples to enhance the labeling process. The complete prompts are provided in Appendix F.2.
@@ 4 Experiments
In this section, we evaluate the performance of representative methods on our GenExam. We also validate the alignment between MLLM-as-a-judge and human evaluation. We use gpt-5-2025-08-07 OpenAI (2025a) as the evaluator model and set the reasoning effort as “low”. More experiments are provided in Appendix C, including results categorized by subject knowledge difficulty and level-2 taxonomy and detailed scores of each dimension.
@@ 4.3 Human Alignment
We conduct a human alignment experiment to examine the validity of MLLM-as-a-judge, following the paradigms in  Zhao et al. (2025); Sun et al. (2025). Five experts are asked to score 250 randomly sampled images generated by GPT-Image-1. They first provide an overall rating between 1 and 10, and then follow the automatic evaluation protocol to annotate scoring points, spelling, logical consistency and readability. The average scores among all experts are used to calculate the mean absolute error (MAE) and correlations between MLLM-judge predicted scores and human scores. In Tab. 3(a), the results show a low MAE across the four dimensions, indicating high precision of our evaluation pipeline.
We then use the overall rating to calculate correlations (Kendall’s $\tau$, Spearman’s $\rho$ and Pearson’s $r$) between human scores and four automatic metrics: our relaxed score, our semantic correctness, VQA score Lin et al. (2024), and CLIP score Hessel et al. (2021). In Tab. 3(b), our relaxed score achieves high correlations with human judgments with $p<0.05$, slightly better than using solely semantic correctness without visual plausibility, while CLIP score and VQA score fail to capture the correctness of multidisciplinary images.
4.4 Ablation on the Evaluator Model
@@
We further study the selection of evaluator model by examining different MLLMs, including closed-source models (GPT-5, GPT-5-mini, o3, o4-mini, GPT-4.1, GPT-4o OpenAI (2025a; c)) and open-source models (Qwen3-VL Bai et al. (2025b), InternVL3.5 Wang et al. (2025a), Intern-S1, Intern-S1-mini Bai et al. (2025a)). On the same manually evaluated samples as in Tab. 3, we run each of the models as evaluator and calculate the accuracy of answering scoring points, MAE of visual plausibility, and three correlation metrics (Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$).
As shown in Tab. 4, the latest model GPT-5 with low reasoning effort achieves the best performance with acceptable cost and time, while using medium or high reasoning effort leads to a potential decrease in performance and higher costs. In addition, using ground truth images as reference help to improve the evaluation precision. Early closed-source models have relatively low performance, especially those without reasoning abilities.
For open-source models, Qwen3-VL demonstrates relative strong performance, but still lags behind GPT-5. Other models have a larger gap between closed-source models, where some scoring point accuracy is close to random guess (50%) and some correlations are even of no statistical significance, suggesting these evaluators are not applicable. This is because multidisciplinary images are still difficult for current multimodal understanding models Yue et al. (2024b). Therefore, it is necessary to use advanced MLLMs to evaluate the generated images for higher precision and robustness.
4.5 Qualitative Analysis
Examples of generated images are shown in Fig. 7. Taking GPT-Image-1 as an example, common error types include lack of knowledge (e.g. keys in the Circle of Semitones), incorrect reasoning (e.g. failing to infer the coordinate of the intersection point), and graph-drawing errors of logical inconsistency (e.g. the point (1,0)(-1,0) appears to be on the y-axis), highlighting the necessity to integrate multimodal understanding, reasoning and generation. Some models show other graph-drawing errors of spelling (e.g. “Transpostion” for Qwen-Image) or readability (e.g. occluded “Music Theory” for HiDream-I1-Full), which is a key capability in multidisciplinary text-to-image exams.
These examples align with our quantitative results: while top closed-source models can partially integrate basic knowledge into generation, they still struggle with precise, domain-aware execution; open-source models, on the contrary, often fail at basic aspects like generating visual plausible images. This underscores the need for future models to prioritize multidisciplinary knowledge integration, rigorous reasoning, and fine-grained visual coherence. Such failures also demonstrate the value of GenExam ’s fine-grained scoring criteria, as they capture flaws that broad metrics like photorealism or general aesthetics would miss. More qualitative results are provided in Appendix G.
5 Conclusion
Ethics Statement
All images in our benchmark are constructed from publicly available resources and existing datasets. We will release the benchmark with clear usage guidelines to prevent misuse for generating misleading content. We have checked for potential biases across all the subjects, aiming for balanced knowledge coverage without favoring specific demographics or disciplines.
Appendix A More Statistics
(a) Biology
(b) Chemistry
(c) Computer Science
(d) Economics
(e) Engineering
(f) Geography
(g) History
(h) Math
(i) Music
(j) Physics
Figure 10: Word clouds of prompts in each subject.
Appendix B Automatic Data Filtering
Appendix C Additional Experiments
In Tab. 5, we show the results on GenExam-Mini, which are consistent with the full dataset. GenExam-Mini can be considered as a valid subset for efficient and affordable validation.
@@ Appendix C Additional Experiments
The results on each level-2 taxonomy is provided in Tab. 8. To highlight the difference between models, and since some models can only obtain a 0% strict score, we only report relaxed scores.
@@
Model Strict Relaxed